OpenStudy (anonymous):

if the period of a clock pendulum on another planet is four times its period on earth, what is the gravity acceleration on the planet? Then calculate the pendulum on the planets moon where g=1.6m/s/s. How would this period change?

OpenStudy (anonymous):

\[T_0=\sqrt{\frac{l}{g}}\] \[(T_0)^2 g=l\] \[ g=\frac{l}{(T_0)^2}\] \[ g=\frac{l}{(4T_0)^2}\] \[ g=\frac{l}{16(T_0)^2}\] so we can see that

OpenStudy (anonymous):

so we can see that gravity on that planet is 1/16 that of earth

OpenStudy (anonymous):

i would it be 1/16 if it is four times? is that because the acceleration is faster? (sorry just missed a lab and trying to work my way through it alone) I was thinking it would be 4times the acceleration meanding the equation you would use on earth=answer (times 4).

OpenStudy (anonymous):

im very confused, so I am probally extremely wrong

OpenStudy (anonymous):

gravity is x1/16 if period x 4

OpenStudy (anonymous):

if gravity(acceleration) is bigger, pendulum will moves faster thus shorter period

OpenStudy (anonymous):

okay concept makes sense now.

OpenStudy (anonymous):

so let me think..the moon would be 1.6 of earth. so less?

OpenStudy (anonymous):

no gravity on the mooon is 1.6

OpenStudy (anonymous):

arghh! so concept not make sense haha

OpenStudy (anonymous):

\[T_0=\sqrt{\frac{l}{g}}\] \[1.6/9.8=.16\] \[G_{earth}*.16=G_{moon}\] \[T_0=\sqrt{\frac{l}{g* .16}}\]

OpenStudy (anonymous):

so l= 1.6? or am i missing something important here?

OpenStudy (anonymous):

im very sorry, and thanks for your patience with me if this should be simple and I should be grasping it

OpenStudy (anonymous):

forget my last line

OpenStudy (anonymous):

\[T_0=\sqrt{\frac{l}{g_{earth}}}\] to get moon's gravity , we multiply earth's .16 \[\frac{T_0}{\sqrt{.16}}=\sqrt{\frac{l}{g_{earth}*.16}}\]

OpenStudy (anonymous):

okay. awesome, so the answer to my first question would be: 1/16th. I can use the equation you gave above to calculate the planets moon. Now if the moon was earths moon. I would not need the *.16 in your last equation. I would just plug in 1.6 or 2.1 or whatever it is & 9.8.

OpenStudy (anonymous):

*9.8 not &9.8

OpenStudy (anonymous):

.16=1.6/9.8

OpenStudy (anonymous):

This problem wants you to find out what happen period with different gravity;

OpenStudy (anonymous):

i think i think I see that. I have a q few more questions on this lab. Mostly exploratory. I will try to to them myself first. Will you be around for awhile in case I get stuck? haha

OpenStudy (anonymous):

sure, just took physics test last night;

OpenStudy (anonymous):

im drowning in physics ...a family member of mine was sick for about 7 weeks. missed 7 labs. trying to do them all before sunday (sunday is the last day my prof will let me take my lab. final grades are due the next day for all labs). he is giving me room to work, but I have lots of lab data, andputting it where it belongs and remembering what its for is hard...even though its organized physics final wedneday.... so thanks for your help.

OpenStudy (anonymous):

sure man, hopefully I remember my physics right

OpenStudy (anonymous):

looks right

OpenStudy (anonymous):

now it actually asks me to do earths moon, so \[ T=\sqrt{\left(lgearth*1.6 \right)}\]

OpenStudy (anonymous):

for the problem above , I did it for earth's moon

OpenStudy (anonymous):

would the 1/16th still apply?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the .16 i mean

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have this all screwed up then...ha, sorry.

OpenStudy (anonymous):

ill scan in the three lab questions. its not making sense to me.

OpenStudy (anonymous):

yep, I will look at it

OpenStudy (anonymous):

our labs are written by yahoos

OpenStudy (anonymous):

That's the case everywhere

OpenStudy (anonymous):

I think students should write these things. Ours do not even follow the lecture schedule

OpenStudy (anonymous):

I think the way they conduct physics lab is really stupid

OpenStudy (anonymous):

agreed. i think the way physics in general is conducted is rather...odd and I like physics haha

OpenStudy (anonymous):

I like physics too

OpenStudy (anonymous):

will you be taking phys II?

OpenStudy (anonymous):

yes, but i failed calc 1, so no go till next january

OpenStudy (anonymous):

i would withdraw, but they would make me drop my physics, which im passing haha

OpenStudy (anonymous):

cpreq mandatory

OpenStudy (anonymous):

coreq*

OpenStudy (anonymous):

what's your major?

OpenStudy (anonymous):

physics

OpenStudy (anonymous):

oh, you really like physics, then

OpenStudy (anonymous):

im good at this stuff, when not rushed, but I am cramming stuff, brain can't take it anymore, so I am here playing stupid.

OpenStudy (anonymous):

I just need three physics

OpenStudy (anonymous):

although this period stuff has me stumped for some reason, as well as the periodic motion of a pendulum

OpenStudy (anonymous):

what your major?

OpenStudy (anonymous):

electrical engineering

OpenStudy (anonymous):

(waiting on scanner to get done)

OpenStudy (anonymous):

see. I think i would enjoy that too

OpenStudy (anonymous):

i probally should have asked if you can open .pdf first

OpenStudy (anonymous):

it's fine

OpenStudy (anonymous):

what school are you at?

OpenStudy (anonymous):

U of I

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

some obscure univerity that no one has heard of but they made a movie about......haha

OpenStudy (anonymous):

Marshall Univ.

OpenStudy (anonymous):

west virginia

OpenStudy (anonymous):

yelp, west virginia I have only met one prof. there that enjoys his job, the rest are there because no one else would hire them, i think......

OpenStudy (anonymous):

(yeah that was mean)

OpenStudy (anonymous):

I am gonna make my own pdf , to answer your question

OpenStudy (anonymous):

haha, well I appreciate it a whole lot.

OpenStudy (anonymous):

this is first one, explained better

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

my labs are a mess. I am pretty sure I am going to fail this...as much as that sucks i have data missing that shouldnt be, there are two pages like this for everyone that is just review concepts. (like 5 of them). i think the data is all here somewhere but i can't make sense out of what its for. my fault for cramming lab experiments together and I know NO ONE elese in the class i can call for help

OpenStudy (anonymous):

second last is just drawing vectors

OpenStudy (anonymous):

now that is something i can do haha

OpenStudy (anonymous):

went ahead and gave you good answer before i forgot

OpenStudy (anonymous):

plot for last one

OpenStudy (anonymous):

you are all set

OpenStudy (anonymous):

yes, thank you. Now to write my stupid error analysis report...so basicall if we (well some scientist did) measured gravity wrong, then its all wrong, or if L is wrong... i have to turn that into a one page paper.....setup dumb....but I can do that!

OpenStudy (anonymous):

i just feel its all setup horribly

OpenStudy (anonymous):

I hated error analysis, we just used to make up numbers

OpenStudy (anonymous):

we have to type a report full of nonsense

OpenStudy (anonymous):

well man, thanks for your help, i think I am gonna take a break. do you frequent the site often?

OpenStudy (anonymous):

if so I will see you around. I have trillian contact information on my about me on openstudy thanks again!

OpenStudy (anonymous):

lol,yes i do frequent this site